OpenStudy (anonymous):

need help with conversions... such as... Diamond (a form of elemental carbon) has a density of 3.513 g cm^-3. What is the density of diamond in pounds foot^-3?

OpenStudy (australopithecus):

First can you tell me what 1g is in pounds and 1cm^3 is in in foot^3?

OpenStudy (anonymous):

thats the problem i dont know the conversions or where to begin... i used to know them but that knowledge has disappeared

OpenStudy (australopithecus):

use wolfram alpha

OpenStudy (australopithecus):

I'm just lazy lol you can use google to but i will do all the work just one second

OpenStudy (australopithecus):

0.002205lb = 1g 3.531*10^-3foot^3 =1 cm^3 knowing this we set up the ratio \[\frac{0.002205lb}{1g}(\frac{1cm^3}{3.531*10^{-3}foot^3}) = \frac{1*0.002205lb(cm^3)}{1*3.531*10^{-3}(g)foot^3}\]

OpenStudy (australopithecus):

now we multiply the density you have to convert it over the units you don't want will cancel out

OpenStudy (australopithecus):

\[\frac{1*0.002205lb(cm^3)}{1*3.531*10^{3}(g)foot^3} * \frac{3.513g}{1cm^{3}} = \frac{(1*0.002205lb(cm^3))*3.513g}{(1*3.531*10^{3}(g)foot^3)(1cm^{3})}\]

OpenStudy (australopithecus):

the conversion ratio should actually be \[\frac{1*0.002205lb(cm^3)}{1*3.531*10^{3}(g)foot^3}\]

OpenStudy (australopithecus):

you just treat the units like numbers they cancel each other out

OpenStudy (australopithecus):

all exponents should be + I made a mistake by making them negative

OpenStudy (australopithecus):

you are left with \[\frac{0.002205*3.513lb}{3.531*10^{3}foot^3}\]

OpenStudy (australopithecus):

do you follow?

OpenStudy (anonymous):

i believe i am starting to... im trying to look through what you did and make something of it

OpenStudy (australopithecus):

just treat the units like numbers they divide to equal 1 so they don't effect the final answer

OpenStudy (australopithecus):

Just an aside when you see a negative exponent it just means that the number belongs in the denominator The rule: \[x^{-1} = \frac{1}{x}\]

OpenStudy (australopithecus):

\[x^{-2} = \frac{1}{x^2}\] etc...

OpenStudy (anonymous):

i know that about exponents but isn't density written as unit^-3

OpenStudy (australopithecus):

\[ 3.513g(1cm^{-3}) = \frac{3.513g}{1cm^{3}} = \frac{3.513g}{cm^{3}} \]

OpenStudy (australopithecus):

note you can divide any number by 1 and it wont change the numbers value

OpenStudy (anonymous):

so that's where that's from...

OpenStudy (australopithecus):

yes negative units mean that the unit belongs in the denominator

OpenStudy (australopithecus):

so I simply set up a ratio to cancel the units

OpenStudy (australopithecus):

if you flip the ratios you can convert back to where you originally start btw

OpenStudy (australopithecus):

since we assume that we can always divide by 1 we just exclude it from the answer but it is there

OpenStudy (australopithecus):

do you follow?

OpenStudy (australopithecus):

The most important thing to do is look at the units like numbers. They will just cancel out like numbers

OpenStudy (anonymous):

alright thank you

OpenStudy (australopithecus):

I think I made a mistake sorry I have a pretty bad migraine it is making mess up a lot all my senses are like super heightened it is really distracting

OpenStudy (australopithecus):

\[\frac{0.002205∗3.513lb}{3.531∗10^{-3}foot^3}\]